# Graph quandles: Generalized Cayley graphs of racks and right quasigroups

### v3

@@ 6. From racks to marked graphs
In this section, we answer the second question in Problem 1.7. We start by addressing a generalized version of the question for right quasigroups and deduce solutions for racks afterward.
@@ 6.2. Specialization to racks
By considering the third conditions in Theorems 6.16.2, we answer the second question in Problem 1.7 in its original form: All racks are realizable by their full Cayley (di)graphs.
@@ 7.1.1. Generalizations of right quasigroups
Conversely, the magma structure given by $R_{y}(x):=x^{3}-x$ on the rational numbers $\mathbb{Q}$ yields a right-divisible magma that is not right-cancellative.
@@ 7.1.2. Labeled digraphs
Labeled edges $(v,\ell,w)\in E$ can also be denoted by transitions $v\xrightarrow{\ell}w$; for example, see Figures 68.
@@ 7.3. First results
Henceforth, let $\mathcal{Q}$ denote the subclass of $\mathcal{D}$ whose elements are also codeterministic and target-complete. The following answers Problem 1.8 for right-cancellative magmas, right-divisible magmas, and right quasigroups.
@@ 7.4. Specialization to racks
    See Figure 6 for a visualization. Note that we do not assume that $w_{1}\neq w_{2}$.
@@
    Figure 6. The rack condition from Definition 7.13 states that for all subgraphs of the form on the left, there also exists a subgraph of the form on the right, where $x:=R_{\ell_{1}}(w_{2})$.
@@
    Figure 7. The second rack condition from Definition 7.13 states that for all subgraphs of the form on the left and non-label vertices $x\in V\setminus L$, if $R_{\ell}(x)\in L$ is a label, then there exists a subgraph of the form on the right.
@@
    Figure 8. The label-involutory condition from Definition 7.13 states that subgraphs of the forms on the left and the right partition the edge set.
@@ 7.4.1. Result
We answer Problem 1.8 for racks, quandles, and involutory right quasigroups. This also yields answers for involutory racks and kei in the obvious way.
@@ 8. Open questions
First, Problem 1.6 motivates the following.
@@
Add more entries to Table 1.
@@
Finally, Problem 1.8 and recent work applying geometric group theory to quandle theory [metrics] motivate further analogues of Theorem 7.14.

### v4

@@ 6. From racks to marked graphs
In this section, we answer the second question in Problem 1.2. We start by addressing a generalized version of the question for right quasigroups and deduce solutions for racks afterward.
@@ 6.2. Specialization to racks
By considering the third conditions in Theorems 6.16.2, we answer the second question in Problem 1.2 in its original form: All racks are realizable by their full Cayley (di)graphs.
@@ 7.1.1. Generalizations of right quasigroups
Conversely, the magma structure given by $R_{y}(x):=x^{3}-x$ on the set of rational numbers $\mathbb{Q}$ yields a right-divisible magma that is not right-cancellative.
@@ 7.1.2. Labeled digraphs
Labeled edges $(v,\ell,w)\in E$ can also be denoted by transitions $v\xrightarrow{\ell}w$; for example, see Figures 7.17.3.
@@ 7.3. First results
Henceforth, let $\mathcal{Q}$ denote the subclass of $\mathcal{D}$ whose elements are also codeterministic and target-complete. The following answers Problem 1.3 for right-cancellative magmas, right-divisible magmas, and right quasigroups.
@@ 7.4. Specialization to racks
    See Figure 7.1 for a visualization. Note that we do not assume that $w_{1}\neq w_{2}$.
@@
    Figure 7.1. The rack condition from Definition 7.13 states that for all subgraphs of the form on the left, there also exists a subgraph of the form on the right, where $x:=R_{\ell_{1}}(w_{2})$.
@@
    Figure 7.2. The second rack condition from Definition 7.13 states that for all subgraphs of the form on the left and non-label vertices $x\in V\setminus L$, if $R_{\ell}(x)\in L$ is a label, then there exists a subgraph of the form on the right.
@@
    Figure 7.3. The label-involutory condition from Definition 7.13 states that subgraphs of the forms on the left and the right partition the edge set.
@@ 7.4.1. Result
We answer Problem 1.3 for racks, quandles, and involutory right quasigroups. This also yields answers for involutory racks and kei in the obvious way.
@@ 8. Open questions
First, Problem 1.1 motivates the following.
@@
Add more entries to Table 5.1.
@@
[…]